The perimeter of the rectangle is 34 cm. Find the value of x.
One side is 11cm and the other is 3x cm.

Answer by harpazo(655)

34 = 2(11) + 2(3x)
34 = 22 + 6x
34-22 = 6x
12 = 6x
12/6 = x
2 = x